Find the P-value for a test of the claim that more than 50% of adults that follow a particular diet will experience increased energy. Of 100 randomly selected subjects who followed the diet, 47 noticed an increase in their energy level.

Question 4 options:

0.2743

0.2257

0.05486

0.7257
Answer #1

Solution :

This hypothesis test is a right tailed test .  

The null and alternative hypothesis is

H0 : p = 0.5

Ha : p > 0.5

\hat p = x / n = 47 / 100 = 0.47

1 - P0 = 1 - 0.5 = 0.5

Test statistic = z

= \hat p - P0 / [\sqrtP0 * (1 - P0 ) / n]

= 0.47 - 0.5 / [\sqrt(0.5 (0.5)) / 100]

= -0.6

P(z > -0.6) = 1 - P(z < -0.6) = 1 - 0.2743 = 0.7257

P-value = 0.7257